# Biased random walk on dynamical percolation

### v2

@@ 1 Introduction
Fix a locally finite graph $G=(V,E)$ and an initial state $\eta\in\{0,1\}^{E(G)}$ of the edges. We say that an edge $e$ is open at time $t$ if $\eta_{t}(e)=1$, and closed otherwise. For parameters $\mu>0$ and $p\in[0,1]$, we consider the dynamics $(\eta_{t})_{t\geq 0}$ with $\eta_{0}=\eta$, where each edge $e$ in the graph is assigned an independent Poisson process of rate $\mu$. If there is a point of the Poisson process at time $t$, we refresh the state of $e$ in $\eta_{t}$, i.e. we declare $e$ open with probability $p$ and closed with probability $1-p$, independently of all other edges and previous states of $e$.
@@
$\lim_{\lambda\rightarrow 0}v^{\prime}(\lambda)=\sigma^{2},$
@@ 2 Regeneration times for the biased random walk
We fix an enumeration $(e_{i})_{i\in\mathbb{N}}$ of the edges in $E$ according to an arbitrary rule. Then for each edge $e_{i}$, we create an infinite number of copies denoted $e_{i,1},e_{i,2},\dots$. We now define a process $(I_{t})_{t\geq 0}$, where for every $t\geq 0$, $I_{t}$ is a set containing copies of edges that we refer to as the infected set.  [^1] Note that in [19] the infected set can contain also the original edges. Let $I_{0}=\emptyset$. Suppose that for some $t\geq 0$, the Poisson clock associated to the random walk $(X_{t})_{t\geq 0}$ rings and that the walker examines the edge $e_{i}$ for some $i\in\mathbb{N}$. If no copy of $e_{i}$ is contained in $I_{t_{-}}$, we set
@@
For all $p\in(0,1)$, for all $\lambda>0$, and for all $\mu>1$, the increments $(\tau_{i}-\tau_{i-1})_{i\in\mathbb{N}}$ are i.i.d., have exponential tails and satisfy $\mathbb{E}_{\lambda}\!\left[\tau_{1}\right]=e^{1/\mu}$.
@@ 3 Speed and Einstein relation
Recall the sequence of regeneration times $(\tau_{i})_{i\in\mathbb{N}}$ from (2.3). Then, $\mathbb{P}_{\lambda}$-almost surely,
$\lim_{t\rightarrow\infty}\frac{X_{t}}{t}=(v(\lambda),0,\dots,0)=\frac{\mathbb{E}_{\lambda}[X_{\tau_{1}}]}{\mathbb{E}[\tau_{1}]}\,.$ (3.1)
@@
It now follows that the first fraction on the right hand side above converges to 00 as $t\to\infty$ almost surely. Using that $\tau_{k}/k\to\mathbb{E}\!\left[\tau_{1}\right]$ as $k\to\infty$ almost surely, it follows easily that the second fraction on the right hand side above converges to 00 almost surely. Finally using (3.3) we get that the third fraction converges to $\mathbb{E}_{\lambda}\!\left[X_{\tau_{1}}\right]/\mathbb{E}\!\left[\tau_{1}\right]$ almost surely as $t\to\infty$ and this concludes the proof. ∎
@@
The biased random walk $(X_{t})_{t\geq 0}=((X_{t}^{1},\ldots,X_{t}^{d}))_{t\geq 0}$ on dynamical percolation satisfies an invariance principle
@@
Recall from Proposition 3.1 that the speed $v(\lambda)$ is equal to $\mathbb{E}_{\lambda}[X_{\tau_{1}}]/\mathbb{E}\!\left[\tau_{1}\right]$ and $\mathbb{E}\!\left[\tau_{1}\right]$ does not depend on $\lambda$. Define the $\sigma$-field ${\cal F}_{\tau_{1}}$ generated by the evolution of the percolation, the exponential waiting times of the particle and the decisions of the particle up to the stopping time $\tau_{1}$. (Note that for the decisions of the particle one can simply generate an i.i.d. sequence of uniform in $[0,1]$ random variables.) Further, define the smaller $\sigma$-field ${\cal G}_{\tau_{1}}$ generated by the evolution of the percolation and the path of the walk up to time $\tau_{1}$. More precisely, ${\cal G}_{\tau_{1}}:=\sigma\left((X_{t},\eta_{t})_{0\leq t\leq\tau_{1}}\right)$. Consider the Radon-Nikodym derivative $\left.\frac{d\mathbb{P}_{\lambda}}{d\mathbb{P}_{0}}\right|_{{\cal F}_{\tau_{1}}}$ of $\mathbb{P}_{\lambda}$ with respect to $\mathbb{P}_{0}$ on ${\cal F}_{\tau_{1}}$. For every $t\geq 0$ we let $R(t)$, respectively $L(t)$, be the number of steps to the right, respectively to the left, that $X^{1}$ performed by time $t$ and $\mathcal{U}(t)$ be the number of steps that were carried out by the walker until time $t$. In the same way, let $R_{a}(t)$, respectively $L_{a}(t)$, be the number of attempted jumps in the $\textup{e}_{1}$ direction, respectively in the $-\textup{e}_{1}$ direction, of $X^{1}$ until time $t$ and recall that $\mathcal{U}_{a}(t)$ is the total number of attempted jumps of $X$ until time $t$. We then have
@@
$\begin{split}&\left.\frac{d\mathbb{P}_{\lambda}}{d\mathbb{P}_{0}}\right|_{{\cal F}_{\tau_{1}}}=e^{\lambda(R_{a}(\tau_{1})-L_{a}(\tau_{1}))}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(\tau_{1})}.\end{split}$ (3.10)
@@
Indeed, conditional on the evolution of the percolation, since it is the same for both measures, for the $\lambda$-biased walk, each attempted jump to the right (resp. to the left) has probability $e^{\lambda}/Z_{\lambda}$ (resp. $e^{-\lambda}/Z_{\lambda}$), while each other direction has probability $1/Z_{\lambda}$. For the unbiased walk the respective probabilities are all equal to $1/(2d)$. We can rewrite the Radon-Nikodym derivative from (3.10) as follows
@@
$\begin{split}&\left.\frac{d\mathbb{P}_{\lambda}}{d\mathbb{P}_{0}}\right|_{{\cal F}_{\tau_{1}}}=e^{\lambda(R(\tau_{1})-L(\tau_{1}))}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}(\tau_{1})}\cdot e^{\lambda(R_{a}(\tau_{1})-R(\tau_{1})-(L_{a}(\tau_{1})-L(\tau_{1}))}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(\tau_{1})-\mathcal{U}(\tau_{1})}\,.\end{split}$ (3.11)
@@
The second factor in the product above is the conditional Radon-Nikodym derivative of $\mathbb{P}_{\lambda}$ with respect to $\mathbb{P}_{0}$ conditional on $\mathcal{G}_{\tau_{1}}$. More precisely,
$\frac{d\mathbb{P}_{\lambda}\!\left(\cdot\;\middle|\;\mathcal{G}_{\tau_{1}}\right)}{d\mathbb{P}_{0}\!\left(\cdot\;\middle|\;\mathcal{G}_{\tau_{1}}\right)}=e^{\lambda(R_{a}(\tau_{1})-R(\tau_{1})-(L_{a}(\tau_{1})-L(\tau_{1}))}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(\tau_{1})-\mathcal{U}(\tau_{1})}.$
Indeed, for each attempted but not executed jump to the right (resp. left) the $\mathbb{P}_{\lambda}$ measure gives a contribution of $e^{\lambda}/Z_{\lambda}$ (resp. $e^{-\lambda}/Z_{\lambda}$), while for the other directions it contributes $1/Z_{\lambda}$. Under $\mathbb{P}_{0}$ all the probabilities are equal to $1/(2d)$. Since the expectation of this conditional Radon-Nikodym derivative is equal to $1$, we get in particular,
$\mathbb{E}\left[\left.\left.\frac{d\mathbb{P}_{\lambda}}{d\mathbb{P}_{0}}\right|_{{\cal F}_{\tau_{1}}}\right|{\cal G}_{\tau_{1}}\right]=e^{\lambda(R(\tau_{1})-L(\tau_{1}))}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}(\tau_{1})}\,.$
Setting $R=R(\tau_{1})$ and $L=L(\tau_{1})$ we can write
$\mathbb{E}_{\lambda}[X_{\tau_{1}}^{1}]=\mathbb{E}_{0}\left[X_{\tau_{1}}^{1}\left.\frac{d\mathbb{P}_{\lambda}}{d\mathbb{P}_{0}}\right|_{{\cal F}_{\tau_{1}}}\right]=\mathbb{E}_{0}\left[(R-L)\cdot e^{\lambda(R-L)}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}(\tau_{1})}\right],$
where we used the fact that $(R-L)$ is measurable with respect to ${\cal G}_{\tau_{1}}$. Hence
$f(\lambda)$ $:=\mathbb{E}_{0}\!\left[X_{\tau_{1}}^{1}e^{\lambda(R-L)}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}(\tau_{1})}\right]$
$=\sum_{m\in\mathbb{N}}\sum_{k+\ell\leq m}(k-\ell)e^{\lambda(k-\ell)}\left(\frac{2d}{Z_{\lambda}}\right)^{m}\mathbb{P}_{0}\left((R,L)=(k,\ell),\mathcal{U}(\tau_{1})=m\right),$ (3.12)
@@
where $\mathbb{P}_{0}$ denotes the law of the symmetric simple random walk on dynamical percolation.
@@
$v^{\prime}(\lambda)=\frac{1}{\mathbb{E}\!\left[\tau_{1}\right]}\cdot\left(\mathbb{E}_{\lambda}\!\left[(X_{\tau_{1}}^{1})^{2}\right]-\frac{e^{\lambda}-e^{-\lambda}}{Z_{\lambda}}\cdot\mathbb{E}_{\lambda}\!\left[X_{\tau_{1}}^{1}\cdot\mathcal{U}(\tau_{1})\right]\right).$
In particular, we have
$\lim_{\lambda\to 0}v^{\prime}(\lambda)=\sigma^{2},$
where $\sigma^{2}$ is the variance from Theorem 1.1.
@@
$\begin{split}\lim_{\delta\to 0}&\frac{f(\lambda+\delta)-f(\lambda)}{\delta}\\ &=\sum_{m\in\mathbb{N}}\sum_{k+\ell\leq m}(k-\ell)e^{\lambda(k-\ell)}\left(\frac{2d}{Z_{\lambda}}\right)^{m}\left(k-\ell-m\cdot\frac{Z_{\lambda}^{\prime}}{Z_{\lambda}}\right)\mathbb{P}_{0}((R,L)=(k,\ell),\mathcal{U}(\tau_{1})=m),\end{split}$ (3.13)
where $Z_{\lambda}^{\prime}:=e^{\lambda}-e^{-\lambda}$. Note that the sum appearing above divided by $\mathbb{E}\!\left[\tau_{1}\right]$ is equal to the expression for the derivative given in the statement of the lemma. A direct calculation shows that
$\frac{f(\lambda+\delta)-f(\lambda)}{\delta}=\sum_{m\in\mathbb{N}}\sum_{k+\ell\leq m}(k-\ell)e^{\lambda(k-\ell)}\left(\frac{2d}{Z_{\lambda}}\right)^{m}g(\delta)\cdot\mathbb{P}_{0}((R,L)=(k,\ell),\mathcal{U}(\tau_{1})=m),$
@@
$g(\delta)=\frac{e^{\delta(k-\ell)}Z_{\lambda}^{m}-Z_{\lambda+\delta}^{m}}{\delta Z_{\lambda+\delta}^{m}}.$
@@
$\left|\frac{f(\lambda+\delta)-f(\lambda)}{\delta}\right|$ $\leq\sum_{m\in\mathbb{N}}\sum_{k+\ell\leq m}Cm^{2}e^{\lambda(k-\ell)}\left(\frac{2d}{Z_{\lambda}}\right)^{m}\mathbb{P}_{0}((R,L)=(k,\ell),\mathcal{U}(\tau_{1})=m)$
$=C\cdot\mathbb{E}_{\lambda}\!\left[(\mathcal{U}(\tau_{1}))^{2}\right]<\infty,$
where for the last bound we used Lemma 2.3, since clearly $\mathcal{U}(\tau_{1})\leq\mathcal{U}_{a}(\tau_{1})$ and the distribution of $\mathcal{U}_{a}(\tau_{1})$ is independent of $\lambda$. We can thus apply the dominated convergence theorem which allows us to differentiate the summands in (3) with respect to $\lambda$ to get (3.13). Applying the dominated convergence theorem again we see that all the terms appearing in the expression for $v^{\prime}(\lambda)$ are continuous functions in $\lambda$, and hence this finishes the proof. ∎
@@
To see that the speed $v(\lambda)=v_{\mu,p}(\lambda)$ is strictly positive for all $\mu,\lambda>0$ and $p\in(0,1]$, note that $\mathbb{P}_{0}((R,L)=(k,\ell),\mathcal{U}(\tau_{1})=m)=\mathbb{P}_{0}((R,L)=(\ell,k),\mathcal{U}(\tau_{1})=m)$ for all $(k,\ell)\in\mathbb{Z}^{2}$ by symmetry of the random walk and the environment law. Thus, we can write
$f(\lambda)$ $=\sum_{m\in\mathbb{N}}\sum_{k+\ell\leq m}(k-\ell)e^{\lambda(k-\ell)}\left(\frac{2d}{Z_{\lambda}}\right)^{m}\mathbb{P}_{0}((R,L)=(k,\ell),\mathcal{U}(\tau_{1})=m)$

### v3

@@ 1 Introduction
Fix a locally finite graph $G=(V,E)$ and an initial state $\eta\in\{0,1\}^{E(G)}$ of the edges. For parameters $\mu>0$ and $p\in[0,1]$, we consider the dynamics $(\eta_{t})_{t\geq 0}$ with $\eta_{0}=\eta$, where each edge $e$ in the graph is assigned an independent Poisson process of rate $\mu$. If there is a point of the Poisson process at time $t$, we refresh the state of $e$ in $\eta_{t}$, i.e. we declare $e$ open (and write $\eta_{t}(e)=1$) with probability $p$ and closed (and write $\eta_{t}(e)=0$) with probability $1-p$, independently of all other edges and previous states of $e$. We then say an edge $e$ is open at time $t$ if $\eta_{t}(e)=1$, and closed otherwise.
@@
$\lim_{\lambda\rightarrow 0}v^{\prime}(\lambda)=\sigma^{2},$ (1.3)
@@ 2 Regeneration times for the biased random walk
We fix an enumeration $(e_{i})_{i\in\mathbb{N}}$ of the edges in $E$ according to an arbitrary rule. Then for each edge $e_{i}$, we create an infinite number of copies denoted $e_{i,1},e_{i,2},\dots$. We now define a process $(I_{t})_{t\geq 0}$, where for every $t\geq 0$, $I_{t}$ is a set containing copies of edges that we refer to as the infected set.  [^1] Note that in [19] the infected set can contain also the original edges. Let $I_{0}=\emptyset$. Suppose that for some $t\geq 0$, the Poisson clock associated with the random walk $(X_{t})_{t\geq 0}$ rings and that the walker examines the edge $e_{i}$ for some $i\in\mathbb{N}$. If no copy of $e_{i}$ is contained in $I_{t_{-}}$, we set
@@
For all $p\in(0,1)$, for all $\lambda\geq 0$, and for all $\mu>0$, the increments $(\tau_{i}-\tau_{i-1})_{i\in\mathbb{N}}$ are i.i.d., have exponential tails and satisfy $\mathbb{E}_{\lambda}\!\left[\tau_{1}\right]=e^{1/\mu}$.
@@ 3 Speed and Einstein relation
Recall the sequence of regeneration times $(\tau_{i})_{i\in\mathbb{N}}$ from (2.3). Then, $\mathbb{P}_{\lambda}$-almost surely
$\lim_{t\rightarrow\infty}\frac{X_{t}}{t}=(v(\lambda),0,\dots,0)=\frac{\mathbb{E}_{\lambda}[X_{\tau_{1}}]}{\mathbb{E}[\tau_{1}]}\,,$ (3.1)
and, writing $(X_{t})_{t\geq 0}=(X_{t}^{1},\ldots,X_{t}^{d})_{t\geq 0}$, we also have
$\lim_{t\rightarrow\infty}\mathbb{E}_{\lambda}\left[\frac{X^{1}_{t}}{t}\right]=\frac{\mathbb{E}_{\lambda}[X^{1}_{\tau_{1}}]}{\mathbb{E}[\tau_{1}]}\,.$ (3.2)
@@
It now follows that the first fraction on the right hand side above converges to 00 as $t\to\infty$ almost surely. Using that $\tau_{k}/k\to\mathbb{E}\!\left[\tau_{1}\right]$ as $k\to\infty$ almost surely, it follows easily that the second fraction on the right hand side above converges to 00 almost surely. Finally, using (3.4) we get that the third fraction converges to $\mathbb{E}_{\lambda}\!\left[X_{\tau_{1}}\right]/\mathbb{E}\!\left[\tau_{1}\right]$ almost surely as $t\to\infty$ and this concludes the proof of the almost-sure convergence. To show (3.2), note that $\|X_{t}/t\|_{1}\leq U_{a}(t)/t$ and $(U_{a}(t)/t)_{t\geq 0}$ is bounded in $L^{2}(\mathbb{P}_{\lambda})$, implying that $(\|X_{t}/t\|_{1})_{t\geq 0}$ is bounded in $L^{2}(\mathbb{P}_{\lambda})$ and in particular uniformly integrable. ∎
@@
The first component $(X^{1}_{t})_{t\geq 0}$ of the biased random walk $(X_{t})_{t\geq 0}=(X_{t}^{1},\ldots,X_{t}^{d})_{t\geq 0}$ on dynamical percolation satisfies an invariance principle
@@
Recall from Proposition 3.1 that the speed $v(\lambda)$ is equal to $\mathbb{E}_{\lambda}[X^{1}_{\tau_{1}}]/\mathbb{E}\!\left[\tau_{1}\right]$ and $\mathbb{E}\!\left[\tau_{1}\right]$ does not depend on $\lambda$. For every $t\geq 0$ we let $R(t)$, respectively $L(t)$, be the number of steps to the right, respectively to the left, that $X^{1}$ performed by time $t$ and $\mathcal{U}(t)$ be the number of steps that were carried out by the walker until time $t$. Let $R_{a}(t)$, respectively $L_{a}(t)$, be the number of attempted jumps in the $\textup{e}_{1}$ direction, respectively in the $-\textup{e}_{1}$ direction, of $X^{1}$ until time $t$ and recall that $\mathcal{U}_{a}(t)$ is the total number of attempted jumps of $X$ until time $t$. Note that $\mathcal{U}_{a}(t)-\mathcal{U}(t)$ is the number of jumps that were attempted but not carried out. Writing $R=R(\tau_{1}),L=L(\tau_{1})$ and $R_{a}=R_{a}(\tau_{1}),L_{a}=L_{a}(\tau_{1})$ we will show the following.
Lemma 3.4.
We have
$\mathbb{E}_{\lambda}[X_{\tau_{1}}^{1}]=\mathbb{E}_{0}\left[(R-L)e^{\lambda(R_{a}-L_{a})}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(\tau_{1})}\right]\,.$ (3.11)
Proof.
Define the $\sigma$-field ${\cal F}_{t}$ generated by the evolution of the percolation, the exponential waiting times of the particle, the decisions of the particle up to time $t$ and the Poisson process $(N_{t})_{t\geq 0}$ used in the definition of the regeneration times. By saying ”evolution of the percolation” we include the times of updates of $\eta=(\eta_{s})_{s\geq 0}$ even if the state of the edge does not change. Consider the Radon-Nikodym derivative $\frac{d\mathbb{P}_{\lambda}}{d\mathbb{P}_{0}}\big{|}_{{\cal F}_{t}}$ of $\mathbb{P}_{\lambda}$ with respect to $\mathbb{P}_{0}$ on ${\cal F}_{t}$. We then have
@@
$\begin{split}&M_{t}:=\left.\frac{d\mathbb{P}_{\lambda}}{d\mathbb{P}_{0}}\right|_{{\cal F}_{t}}=e^{\lambda(R_{a}(t)-L_{a}(t))}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(t)}.\end{split}$ (3.12)
@@
Indeed, conditional on the evolution of the percolation, since it is the same for both measures, for the $\lambda$-biased walk, each attempted jump to the right (resp. to the left) has probability $e^{\lambda}/Z_{\lambda}$ (resp. $e^{-\lambda}/Z_{\lambda}$), while each other direction has probability $1/Z_{\lambda}$. For the unbiased walk the respective probabilities are all equal to $1/(2d)$. Note that $\tau_{1}$ is a stopping time with respect to the filtration $({\cal F}_{t})_{t\geq 0}$, and that $(M_{t\wedge\tau_{1}})_{t\geq 0}$ is a martingale with respect to $({\cal F}_{t})_{t\geq 0}$. Since $\mathbb{P}_{\lambda}$ is absolutely continuous with respect to $\mathbb{P}_{0}$ on ${\cal F}_{\tau_{1}}$ (this follows simply from the fact that $\tau_{1}<\infty$ almost surely), $(M_{t\wedge\tau_{1}})_{t\geq 0}$ is a uniformly integrable martingale. Hence,
@@
$\begin{split}&M_{\tau_{1}}=\left.\frac{d\mathbb{P}_{\lambda}}{d\mathbb{P}_{0}}\right|_{{\cal F}_{\tau_{1}}}=e^{\lambda(R_{a}-L_{a})}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(\tau_{1})},\end{split}$ (3.13)
@@
i.e. $M_{\tau_{1}}$ is the Radon-Nikodym derivative of $\mathbb{P}_{\lambda}$ with respect to $\mathbb{P}_{0}$ on ${\cal F}_{\tau_{1}}$. Hence
$\mathbb{E}_{\lambda}[X_{\tau_{1}}^{1}]=\mathbb{E}_{0}\left[X_{\tau_{1}}^{1}e^{\lambda(R_{a}-L_{a})}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(\tau_{1})}\right]=\mathbb{E}_{0}\left[(R-L)e^{\lambda(R_{a}-L_{a})}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(\tau_{1})}\right]\,,$ (3.14)
which gives the result. ∎
Let
$p_{0}(k_{a},\ell_{a},k,\ell,m):=\mathbb{P}_{0}\left((R_{a},L_{a})=(k_{a},\ell_{a}),(R,L)=(k,\ell),\mathcal{U}_{a}(\tau_{1})=m\right),$ (3.15)
@@
where $\mathbb{P}_{0}$ denotes the law of the symmetric simple random walk on dynamical percolation. Then
$f(\lambda)$ $:=\mathbb{E}_{\lambda}[X_{\tau_{1}}^{1}]=\mathbb{E}_{0}\!\left[X_{\tau_{1}}^{1}e^{\lambda(R_{a}-L_{a})}\left(\frac{2d}{Z_{\lambda}}\right)^{\mathcal{U}_{a}(\tau_{1})}\right]$
$=\sum_{m\in\mathbb{N}}\sum_{\begin{subarray}{c}k_{a}+\ell_{a}\leq m\\ k\leq k_{a},\ell\leq\ell_{a}\end{subarray}}(k-\ell)e^{\lambda(k_{a}-\ell_{a})}\left(\frac{2d}{Z_{\lambda}}\right)^{m}p_{0}(k_{a},\ell_{a},k,\ell,m)\,.$ (3.16)
@@
$v^{\prime}(\lambda)=\frac{1}{\mathbb{E}\!\left[\tau_{1}\right]}\cdot\left(\mathbb{E}_{\lambda}\!\left[X_{\tau_{1}}^{1}\left(R_{a}-L_{a}\right)\right]-\frac{e^{\lambda}-e^{-\lambda}}{Z_{\lambda}}\cdot\mathbb{E}_{\lambda}\!\left[X_{\tau_{1}}^{1}\cdot\mathcal{U}_{a}(\tau_{1})\right]\right).$ (3.17)
@@
$\begin{split}\lim_{\delta\to 0}&\frac{f(\lambda+\delta)-f(\lambda)}{\delta}\\ &=\sum_{m\in\mathbb{N}}\sum_{\begin{subarray}{c}k_{a}+\ell_{a}\leq m\\ k\leq k_{a},\ell\leq\ell_{a}\end{subarray}}(k-\ell)e^{\lambda(k_{a}-\ell_{a})}\left(\frac{2d}{Z_{\lambda}}\right)^{m}\left(k_{a}-\ell_{a}-m\cdot\frac{Z_{\lambda}^{\prime}}{Z_{\lambda}}\right)p_{0}(k_{a},\ell_{a},k,\ell,m),\end{split}$ (3.18)
where $Z_{\lambda}^{\prime}:=e^{\lambda}-e^{-\lambda}$. The last term equals
$\mathbb{E}_{\lambda}\!\left[X_{\tau_{1}}^{1}(R_{a}-L_{a})\right]-\frac{e^{\lambda}-e^{-\lambda}}{Z_{\lambda}}\cdot\mathbb{E}_{\lambda}\!\left[X_{\tau_{1}}^{1}\cdot\mathcal{U}_{a}(\tau_{1})\right]\,.$
Note that the sum appearing above divided by $\mathbb{E}\!\left[\tau_{1}\right]$ is equal to the expression for the derivative given in the statement of the lemma.
A direct calculation shows that
$\frac{f(\lambda+\delta)-f(\lambda)}{\delta}=\sum_{m\in\mathbb{N}}\sum_{\begin{subarray}{c}k_{a}+\ell_{a}\leq m\\ k\leq k_{a},\ell\leq\ell_{a}\end{subarray}}(k-\ell)e^{\lambda(k_{a}-\ell_{a})}\left(\frac{2d}{Z_{\lambda}}\right)^{m}g(\delta)\cdot p_{0}(k_{a},\ell_{a},k,\ell,m),$
@@
$g(\delta)=\frac{e^{\delta(k_{a}-\ell_{a})}Z_{\lambda}^{m}-Z_{\lambda+\delta}^{m}}{\delta Z_{\lambda+\delta}^{m}}.$
@@
$\left|\frac{f(\lambda+\delta)-f(\lambda)}{\delta}\right|$ $\leq\sum_{m\in\mathbb{N}}\sum_{\begin{subarray}{c}k_{a}+\ell_{a}\leq m\\ k\leq k_{a},\ell\leq\ell_{a}\end{subarray}}Cm^{2}e^{\lambda(k_{a}-\ell_{a})}\left(\frac{2d}{Z_{\lambda}}\right)^{m}p_{0}(k_{a},\ell_{a},k,\ell,m)$
$=C\cdot\mathbb{E}_{\lambda}\!\left[(\mathcal{U}_{a}(\tau_{1}))^{2}\right]<\infty,$
where for the last bound we used Lemma 2.3, since the distribution of $\mathcal{U}_{a}(\tau_{1})$ does not depend on $\lambda$. We can thus apply the dominated convergence theorem which allows us to differentiate the summands in (3) with respect to $\lambda$ to get (3.18). Applying the dominated convergence theorem again we see that all the terms appearing in the expression for $v^{\prime}(\lambda)$ are continuous functions in $\lambda$, and this finishes the proof of (3.17). ∎
Proposition 3.6.
We have
$\lim_{\lambda\to 0}v^{\prime}(\lambda)=\sigma^{2},$
where $\sigma^{2}=\mathbb{E}_{0}\!\left[(X_{\tau_{1}}^{1})^{2}\right]/\mathbb{E}\!\left[\tau_{1}\right]$ is the variance from Theorem 1.1, see (1.3).
Proof.
We see, taking $\lambda\to 0$ in (3.17) the statement follows if we show that
$\mathbb{E}_{0}\!\left[X_{\tau_{1}}^{1}\left(R_{a}-L_{a}\right)\right]=\mathbb{E}_{0}\!\left[X_{\tau_{1}}^{2}\right]\,.$ (3.20)
Turning to the proof of (3.20), we first note that (3.20) is equivalent to
$\mathbb{E}_{0}\!\left[(R-L)\left(R_{\rm supp}-L_{\rm supp}\right)\right]=0,$ (3.21)
where we write $R_{\rm supp}=R_{a}-R$ and $L_{\rm supp}=L_{a}-L$ for the jumps to the right or left respectively that were not carried out (the subscript stands for ”suppressed jumps”). Similarly as in the proof of Proposition 3.1, one can show that, with $R_{\rm supp}(t)=R_{a}(t)-R(t)$, $L_{\rm supp}(t)=L_{a}(t)-L(t)$,
$\lim\limits_{t\to\infty}\frac{1}{t}\mathbb{E}_{0}\!\left[(R(t)-L(t))\left(R_{\rm supp}(t)-L_{\rm supp}(t)\right)\right]=\frac{1}{\mathbb{E}[\tau_{1}]}\mathbb{E}_{0}\!\left[(R-L)\left(R_{\rm supp}-L_{\rm supp}\right)\right]\,.$ (3.22)
Hence, it suffices to prove that for all $t>0$,
$\mathbb{E}_{0}\!\left[(R(t)-L(t))\left(R_{\rm supp}(t)-L_{\rm supp}(t)\right)\right]=0\,.$ (3.23)
We will provide a measure-preserving transformation such that $R(t)-L(t)$ is an antisymmetric function but $R_{\rm supp}(t)-L_{\rm supp}(t)$ is a symmetric function under this transformation, implying that (3.23) holds true. Indeed, the law of the environment and the walk under $\mathbb{P}_{0}$ is invariant under time-reversal. More precisely, fix $t>0$ and let $\tilde{\eta}_{s}=\eta_{t-s}$, $0\leq s\leq t$ and $\tilde{X}_{s}=X_{t-s}-X_{t}$, $0\leq s\leq t$. Consider all the times and decisions of the particle in the path $(X_{s})_{0\leq s\leq t}$ concerning jumps to the right or to the left that were suppressed. Call these times $r_{1},r_{2},\ldots,r_{R_{\rm supp}(t)}$ (for suppressed jumps to the right) and $\ell_{1},\ell_{2},\ldots,\ell_{L_{\rm supp}(t)}$ (for suppressed jumps to the left) respectively. Note that
$\left((\eta_{s})_{0\leq s\leq t},(X_{s})_{0\leq s\leq t},(r_{1},r_{2},\ldots,r_{R_{\rm supp}(t)}),(\ell_{1},\ell_{2},\ldots,\ell_{L_{\rm supp}(t)})\right)$
has the same law under $\mathbb{P}_{0}$ as
$\left((\tilde{\eta}_{s})_{0\leq s\leq t},(\tilde{X}_{s})_{0\leq s\leq t},(t-r_{1},\ldots,t-r_{R_{\rm supp}(t)}),(t-\ell_{1},\ldots,t-\ell_{L_{\rm supp}(t)})\right)\,.$
Since $R(t)-L(t)=X_{t}=-\tilde{X}_{t}$ and $R_{\rm supp}(t)$ and $L_{\rm supp}(t)$ are the same for both processes, see Figure 4, we conclude that (3.23) holds true. ∎
1111222233334444555566667777888899991010101010101010999988887777666655554444333322221111
Figure 4: Visualization of the trajectory of a random walk in dynamical percolation and its time-reversal. The black dot marks the origin $(0,0)$, while the numbers refer to the order in which jumps are attempted (and performed) in the two processes. A line is dotted when a jump was attempted, but suppressed.
@@
We have, using again the time-reversal argument in the proof of Proposition 3.6 for the third equality